# Efficient Transformer Encoders for Mask2Former-style models

### v1

@@ 4 Experiments
Training settings.
The experimental setup closely mirrors that of M2F [5], with all model configurations and training specifics following identical procedures. We use Detectron2 [34] and PyTorch[25] for our implementation. For the stochastic depth training phase (Step A), we initialize weights as provided by M2F and subsequently train 50 epochs for the COCO dataset and 90k iterations for Cityscapes, with a batch size of 16. For the training of the gating network (Step C), we perform 2 epochs of training on the COCO dataset and 20k iterations on the Cityscapes dataset, employing the Adam optimizer [18]. The adaptation factor $\beta$ in the utility function, as discussed in Sec. 3.3, is set to 0.0005 for COCO and 0.003 for Cityscapes, unless otherwise specified. Distributed training is performed using 8 A6000 GPUs. On the COCO dataset, the training time of Step A is 280 GPU hours, Step B is 17 GPU hours, and Step C 7.2 GPU hours. Similarly for Cityscapes dataset, the training time of Step A is 45 GPU hours, Step B is 1 GPU hours, and Step C is 7.2 GPU hours. In Step A, we use identical settings as M2F for the loss between the predicted segment and ground truth segment, i.e., $\mathcal{L}_{k}$. The weight $\lambda_{\text{mask}}$ is fixed at 5.0, while $\lambda_{\text{class}}$ is set to 2.0 for all classes, except 0.1 for the “no object” class.
@@ 4.1 Main Results
Table 1: COCO evaluation. Our method, Task-specific architectures
@@
Model PQ mIoUp APp Total Tx. Enc.
@@
Table 2: Cityscapes evaluation. Our method, Task-specific architectures
@@
Model PQ mIoUp APp Total Tx. Enc.
@@
In Table 2 and Table 2, we compare ECO-M2F with our baseline prior works on the validation set of COCO and Cityscapes dataset, respectively. In Table 2, we observe that ECO-M2F effectively reduces computational costs while upholding performance levels in comparison to M2F [5] using both SWIN-T [22] and Res50 [13] backbones. Additionally, ECO-M2F can be seamlessly integrated into efficient encoder designs, such as Lite-M2F [5][20], further reducing GLOPs by approximately 12.6%. With Res50 as the backbone, MF [6], YOSO [15], and RAP-SAM [40] exhibit inferior performance compared to ECO-M2F. Although ReMax [26] demonstrates competitive accuracy, its focus on specialized panoptic segmentation models limits its applicability. Our work, however, aims for a broader impact by creating efficient segmentation architectures that can be used for various segmentation tasks. We make similar observations on the Cityscapes dataset as presented in Table 2.
@@ 4.2 Ablation Studies
Balancing computational cost and performance.
Within ECO-M2F, the parameter $\beta$ serves as an adaptation factor that governs the trade-off between computational cost and performance. Its value, however, is contingent upon the backbone and dataset characteristics. This dependency arises due to the disparate ranges of GFLOPs and segmentation quality (represented by PQ) which are a function of the architecture components and the training data distribution. Fig. 5 illustrates ECO-M2F with different values of $\beta$ in Step C using the exact same weights for the parent model from Step A during inference. In comparison to the M2Fi model (where each is trained standalone with $i$ layers), adjusting the value of $\beta$ provides a trade-off between GFLOPs and PQ.
Table 3: Impact of $\beta$. As the value of $\beta$ increases, the model places greater emphasis on reducing GFLOPs over performance both in COCO and Cityscapes datasets. Baseline is M2F [5]. (Backbone: SWIN-T)
@@
Data $\beta$ PQ mIoUp APp Total Tx. Enc.
@@
Table 4: Impact of $K$. The baseline M2F [5] refers to the complete M2F model trained with 6/5/4 layers for 50 epochs. ECO-M2F pertains to the training of only the gating network for 2 epochs by setting $K=6,5,4$ respectively. (Dataset: COCO; Backbone: SWIN-T)
Performance ($\uparrow$) GFLOPs ($\downarrow$)
$K$ Model PQ mIOUp APp Total Tx. Enc.
6 M2F [5] 52.03 62.49 42.18 235.57 121.69
ECO-M2F 52.06 62.76 41.51 202.39 88.47
5 M2F[5] 51.61 61.93 41.55 221.95 108.07
ECO-M2F 52.26 62.67 41.56 208.82 94.59
4 M2F[5] 51.38 62.30 41.11 208.33 94.45
ECO-M2F 52.20 62.58 41.55 202.47 88.65
  • Impact of adaptation factor $\beta$. We analyze the impact of $\beta$ on ECO-M2F and present our analysis in Table 4. As expected, a smaller $\beta$ prioritizes segmentation quality over computations resulting in superior performance. Conversely, a larger $\beta$ signifies a greater emphasis on GFLOPs. This results in a slight sacrifice in PQ leading to a significant reduction in GFLOPs.
  • Impact of total encoder layers $K$ in the parent architecture. In situations where computational resources are limited, we present an approach using ECO-M2F to create a scaled-down version of the parent model. As illustrated in Tab. 4, we analyze the impact of $K$ set to 5 and 4 (in place of the default value of 6). We set $\beta=0.0005$ for all the models. It can be observed that ECO-M2F not only reduces the computational load but also preserves the performance of the parent model. Importantly, we only execute Step C for each configuration which involves training the gating network.
Table 5: Stochastic Depth (SD) training. Here, all models (w/ 6 encoder layers) deterministically exit at the marked layer at inference. “USD”: Unweighted; “WSD”: Weighted. (Baseline: M2F w/ SWIN-T)
PQ ($\uparrow$)
Model 2 3 4 5 6
Dataset: Cityscapes
Baseline 03.21 14.37 26.66 42.50 64.00
w/ USD 59.96 61.85 63.18 62.98 63.73
w/ WSD 60.71 62.14 62.94 63.89 64.60
Dataset: COCO
Baseline 10.16 17.02 23.43 33.63 51.71
w/ USD 49.40 50.25 50.49 50.51 50.44
w/ WSD 50.70 51.76 52.30 52.39 52.48
Figure 5: M2Fi is trained w/ total $i$ layers. $\ell_{i}$ is result of same model from our Step A. Values denote $\beta$. Dataset: COCO.
[…]
Performance ($\uparrow$) GFLOPs ($\downarrow$)
Bakbone Model PQ mIOUp APp Total Tx. Enc.
Dataset: COCO
SWIN-T M2F [5] 52.03 62.49 42.18 235.57 121.69
ECO-M2F 52.06 62.76 41.51 202.39 88.47
SWIN-S M2F [5] 54.63 64.24 44.69 316.50 121.69
ECO-M2F 54.76 64.46 44.48 275.96 81.05
SWIN-B M2F 56.40 67.09 46.29 470.98 122.56
ECO-M2F 56.49 66.56 46.40 425.17 76.50
Dataset: Cityscapes
SWIN-T M2F [5] 64.00 80.77 39.26 537.85 281.13
ECO-M2F 64.18 80.49 39.64 507.51 250.80
SWIN-S M2F [5] 64.84 81.76 40.73 724.29 281.13
ECO-M2F 65.12 81.64 41.17 665.38 222.22
SWIN-B M2F [5] 66.12 82.70 42.84 1051.19 283.14
ECO-M2F 65.44 82.05 40.51 984.73 216.69
Impact of backbone features.
We analyze the impact of backbone features on ECO-M2F in Tab. 6 and Tab. 8, since our gating network mechanism takes these features as input.
  • Size variations. In Table 6, we evaluate ECO-M2F with backbones of different sizes. Specifically, we observe the performance with SWIN-Tiny (T), SWIN-Small (S), and SWIN-Base (B) architectures [22]. We can observe the adaptability of ECO-M2F in delivering robust performance efficiency across a range of these backbone sizes. Furthermore, this versatility of ECO-M2F extends to Lite-M2F meta-architecture as well (see supplementary material).
  • Pre-trained weight variations. We explore how different pre-training strategies for the backbones affect the performance of ECO-M2F in Table 8. We initialize the backbone weights using both supervised learning (SL) and self-supervised learning (SSL) techniques on the ImageNet-1K dataset [8]. For SSL pre-training on the SWIN-T [22] backbone, we utilize MoBY [35]. For SSL pre-training on the Res50 [13] backbone, we use DINO [3]. With MoBY weights on the SWIN-T backbone, ECO-M2F maintains 99.7% of the PQ of M2F while reducing GFLOPs in the transformer encoder by 26.34%. With DINO weights on the Res50 backbone, we observe a reduction of 29.79% in GFLOPs in the transformer encoder, alongside a slight improvement in PQ.
Table 7: Impact on DETR. We extend our approach to DETR [2] for object detection tasks as ECO-DETR. (Dataset: COCO; Backbone: Res50)
Performance ($\uparrow$) GFLOPs ($\downarrow$)
Model AP AP50 APS APM APL Total Tx. Enc.
DETR [2] 42.0 62.4 20.5 45.8 61.1 83.59 9.92
ECO-DETR($\beta=0.0001$) 41.9 62.2 20.8 45.8 60.4 81.87 6.41
ECO-DETR($\beta=0.001$) 40.9 61.5 20.2 44.5 59.1 79.92 4.52
ECO-DETR($\beta=0.01$) 40.2 61.0 18.7 43.6 58.8 79.33 3.94
Table 8: Impact of backbone weights. ECO-M2F applies to backbone weights obtained through self-supervised pre-training as well. (Dataset: COCO)
Performance ($\uparrow$) GFLOPs ($\downarrow$)
Bakbone Model PQ mIOUp APp Total Tx. Enc.
SWIN-T(MoBY[36]) M2F[5] 51.64 62.42 41.93 235.57 121.69
ECO-M2F 51.48 62.49 41.27 203.56 89.64
Res50(DINO[3]) M2F[5] 51.57 61.65 41.24 229.41 126.05
ECO-M2F 51.89 61.67 41.30 191.97 88.50
Performance in object detection.
To demonstrate the applicability of our method beyond segmentation, we extend our proposed three step recipe for object detection task using DETR [2] for object detection task and name it ECO-DETR. In particular, we vary the values of the adaptation factor $\beta$ to analyze performance-computation trade-offs of ECO-DETR in Table 8. Clearly, the resultant architecture maintains the performance but helps in reducing the computations of the encoder (e.g., it achieves a 35.38% reduction in GFLOPs in the transformer encoder without significantly impacting performance).
4.3 Qualitative Comparisons
Figure 6: Qualitative visualizations. We illustrate few examples of predicted segmentation maps from M2F [5] (middle column) and ECO-M2F (last column). Top two rows are from the COCO dataset, whereas bottom two rows are from the Cityscapes dataset. (Backbone: Swin-T) Zoom-in for best view.
We present a few examples of predicted segmentation maps in Fig. 6 with SWIN-T [22] backbone. Compared to the parent architecture, ECO-M2F consistently shows strong performance while self-selecting the encoder layers based on the input examples, both in everyday scenes (on COCO dataset) as well as intricate traffic scenes (on Cityscapes dataset).
5 Conclusions
In this paper, we propose an efficient transformer encoder design ECO-M2F for the Mask2Former-style frameworks. ECO-M2F provides a three-step training recipe that can be used to customize the transformer encoder on the fly given the input image. The first step involves training the parent model to be dynamic by allowing stochastic depths at the transformer encoder. The second step involves creating a derived dataset from the training dataset which contains a pair of image and layer number that provides the highest segmentation quality. Finally, the third step involves training a gating network, whose function is to decide the number of layers to be used given the input image. Extensive experiments demonstrate that ECO-M2F achieves significantly reduced computational complexity compared to established methods while maintaining competitive performance in universal segmentation. Our results highlight ECO-M2F’s ability to dynamically trade-off between performance and efficiency as per requirements, showcasing its adaptability across diverse architectural configurations, and can be applied to models for object detection tasks.
Limitations.
Efficient Transformer Encoders for Mask2Former-style models
(Supplementary Material)
A Additional Experiments
Impact of backbone size on Lite-M2F.
We apply ECO-M2F on Lite-M2F using various backbone sizes, including SWIN-Tiny (T), SWIN-Small (S), and SWIN-Base (B) architectures [22]. Lite-M2F is a specific variant based on Lite-DETR [20]. We used the configuration named “Lite-DETR H3L1-(6+1)$\times$1” given its strong performance in detection relative to the computations required. However, we adjust this configuration to (5+1) when applying our approach to Lite-M2F. Further, we use their without the key-aware deformable attention [20] proposed in their paper. This adjustment is necessary because Lite-M2F actually has 6 encoder layers, and the original configuration might introduce an additional layer that isn’t present in the model. Following this, we identify layers 2 to 5 as potential exits, followed by the last layer, layer 6 in the transformer encoder. We retain layer 6 and do not consider it as a feasible exit point as it leverages features from all scales provided by the backbone, making it essential to the model’s functionality. As shown in Tab. T2, we observe that ECO-M2F effectively reduces computational cost while maintaining performance across Lite-M2F variants, which underscores the versatility and robustness of ECO-M2F across different model architectures and sizes.
Impact of target and loss settings for gating network training.
We investigate various target and loss settings during the training of the gating network. Specifically, we compare the approach detailed in the main paper, using one-hot target and cross-entropy loss (referred to as “hard-CE” in Tab. T2), with three alternative methods that do not involve setting a specific target exit for each image.
@@ A Additional Experiments
The analysis in Tab. T2 is conducted using the SWIN-T [22] backbone on the COCO dataset. We observe that “hard-CE” yields the most favorable results. As a result, we use this approach consistently in the main paper.
Table T1: Impact of backbone size on Lite-M2F. Our Lite-ECO-M2F maintains the performance of Lite-M2F while reducing GFLOPs for different datasets and for different backbones.
@@
Bakbone Model PQ mIOUp APp Total Tx. Enc.
Dataset: Cityscapes
SWIN-T Lite-M2F [20] 62.29 79.43 36.57 428.71 172.00
Lite-ECO-M2F 62.64 79.99 36.52 412.88 156.17
SWIN-S Lite-M2F [20] 63.54 79.74 39.12 615.15 171.99
Lite-ECO-M2F 63.32 80.21 37.91 588.82 145.66
SWIN-B Lite-M2F [20] 64.48 82.34 39.21 942.05 174.01
Lite-ECO-M2F 64.66 81.40 39.52 921.15 153.11
Dataset: COCO
SWIN-T Lite-M2F [20] 52.70 63.08 41.10 193.79 79.78
Lite-ECO-M2F 52.84 63.23 42.18 178.43 64.42
SWIN-S Lite-M2F [20] 54.30 64.81 43.94 269.26 74.45
Lite-ECO-M2F 54.47 64.14 43.55 258.00 63.96
Table T2: Impact of target and loss in gating network training. We use “hard-CE” loss for training our gating network in the main paper. (Backbone: SWIN-T; Dataset: COCO)
Performance ($\uparrow$) GFLOPs ($\downarrow$)
Method PQ mIOUp APp Total Tx. Enc.
@@
B Additional Qualitative Results
We provide additional examples of predicted segmentation maps in Fig. F1.
Figure F1: Qualitative visualizations. We provide additional examples of predicted segmentation maps from M2F [5] (middle column) and ECO-M2F (last column). Top two rows are from the COCO dataset, whereas bottom two rows are from the Cityscapes dataset. Please zoom in for a clearer view of the details. (Backbone: SWIN-T)

### v2

@@ 3.5 Experiments Settings
Training Settings
The experimental setup closely mirrors that of M2F [5], with all model configurations and training specifics following identical procedures. We use Detectron2 [34] and PyTorch[25] for our implementation. For the stochastic depth training phase (Step A), we initialize weights as provided by M2F and subsequently train 50 epochs for the COCO dataset and 90k iterations for Cityscapes, with a batch size of 16. For the training of the gating network (Step C), we perform 2 epochs of training on the COCO dataset and 20k iterations on the Cityscapes dataset, employing the Adam optimizer [18]. The adaptation factor $\beta$ in the utility function, as discussed in Sec. 3, is set to 0.0005 for COCO and 0.003 for Cityscapes, unless otherwise specified. Distributed training is performed using 8 A6000 GPUs. On the COCO dataset, the training time of Step A is 280 GPU hours, Step B is 17 GPU hours, and Step C 7.2 GPU hours. Similarly for Cityscapes dataset, the training time of Step A is 45 GPU hours, Step B is 1 GPU hours, and Step C is 7.2 GPU hours. In Step A, we use identical settings as M2F for the loss between the predicted segment and ground truth segment, i.e., $\mathcal{L}_{k}$. The weight $\lambda_{\text{mask}}$ is fixed at 5.0, while $\lambda_{\text{class}}$ is set to 2.0 for all classes, except 0.1 for the “no object” class.
@@ 3.6 Main Results
Model PQ mIoUp APp Total Encoder
@@
Table 1: Evaluation on COCO dataset.. Our method achieves competitive performance with notable reductions in GFLOPs and can be combined with other efficient encoder designs to further enhance overall efficiency. Task-specific architectures
@@
Model PQ mIoUp APp Total Encoder
@@
Table 2: Evaluation on Cityscapes dataset. Our method balances strong performance with lower GFLOPs, making it more efficient than other models that excel in only one aspect. Task-specific architectures
Method $\beta$ PQ mIoUp APp GFLOPs
M2F - 39.71 46.09 26.49 139.81
ECO-M2F 0.0005 38.79 44.71 24.81 127.21
ECO-M2F 0.0010 38.80 44.22 24.78 118.23
ECO-M2F 0.0020 38.63 44.11 24.59 116.95
Table 3: Evaluation on ADE20K dataset. The performance of ECO-M2F on the ADE20K dataset highlights its capability to adaptively balance efficiency and effectiveness. (backbone: Res50)
As shown in Fig. 2 (b), ECO-M2F accuracy-efficiency trade-off can be better than the baseline models. Detailed comparisons are provided in Tab. 1 and Tab. 2, where we benchmark ECO-M2F against baseline prior works on the validation set of COCO and Cityscapes datasets, respectively. In Tab. 1, we observe that ECO-M2F effectively reduces computational costs while upholding performance levels in comparison to M2F [5] using both SWIN-T [22] and Res50 [13] backbones. Additionally, ECO-M2F can be seamlessly integrated into efficient encoder designs, such as Lite-M2F [5] [20], further reducing GLOPs by approximately 12.6%. With Res50 as the backbone, MF [6], YOSO [15], and RAP-SAM [38] underperform compared to ECO-M2F. While ReMax [26] shows competitive accuracy, its specialization in panoptic segmentation limits its general applicability. Our work, however, aims for a broader impact by creating efficient segmentation architectures that can be used for various segmentation tasks and enhance the parent architecture without the size constraints of ReMax. We observe similar results on the Cityscapes dataset (see Tab. 2) and on the ADE20K dataset [43] (as shown in Tab. 3).
@@ 3.7 Ablation Studies
Impact of Adaptation Factor $\beta$
We analyze the impact of $\beta$ on ECO-M2F and present our analysis in Tab. 4. As expected, a smaller $\beta$ prioritizes segmentation quality over computations, resulting in superior performance. Conversely, a larger $\beta$ signifies a greater emphasis on GFLOPs. This results in a slight sacrifice in PQ leading to a significant reduction in GFLOPs.
@@
Data $\beta$ PQ mIoUp APp Total Encoder
@@
Table 4: Impact of the adaptation factor $\beta$. As the value of $\beta$ increases, the model places greater emphasis on reducing GFLOPs over performance both in COCO and Cityscapes datasets. Baseline here is M2F [5]. (Backbone: SWIN-T)
ECO-M2F vs. M2F w/ Weighted Stochastic Depth (WSD)
We present a comparison between M2F trained with WSD and ECO-M2F in Fig. 5. The blue dots denote the results when M2F w/ WSD exits always at $i$th layer ($i=2,\cdots,6$). The red “x”s indicate the performance of our approach with the introduction of the Gating module, which varies with changes in $\beta$. The results clearly demonstrate that ECO-M2F effectively self-selects encoder layers for compute reduction while preserving performance, illustrating its efficiency in optimizing the trade-off between computation and accuracy.
Figure 5: ECO-M2F outperforms M2F with WSD by effectively reducing computation without compromising performance, as shown by the varying results with the Gating module across different $\beta$ values. (Dataset: Cityscapes)
Impact of Target and Loss Settings for Gating Network Training
We investigate various target and loss settings during the training of the gating network. Specifically, we compare the approach detailed in Sec. 3.3, using one-hot target and cross-entropy loss (referred to as “hard-CE” in Tab. 5), with three alternative methods that do not involve setting a specific target exit for each image.
@@
The analysis in Tab. 5 is conducted using the SWIN-T [22] backbone on the COCO dataset. We observe that “hard-CE” yields the most favorable results. As a result, all of the other results use this approach.
@@
Method PQ mIOUp APp Total Encoder
@@
Table 5: Impact of target and loss in gating network training. We use “hard-CE” loss for training our gating network in all of the other results. (Backbone: SWIN-T; Dataset: COCO)
Batch Inference and FPS Comparison
ECO-M2F maintains efficient batch processing by utilizing a shared decoder across all encoder exit points, ensuring that throughput is not compromised. As shown in Fig. 6, our method significantly improves the frames per second (FPS) across various batch sizes, demonstrating a clear advantage in processing speed.
Figure 6: FPS vs. batch size. ECO-M2F consistently achieves higher FPS than M2F [5] across all batch sizes. (Dataset: Cityscapes; Backbone: Res50)
4 Conclusions
In this paper, we propose an efficient transformer encoder design ECO-M2F for the Mask2Former-style frameworks. ECO-M2Fprovides a three-step training recipe that can be used to customize the transformer encoder on the fly, given the input image. The first step involves training the parent model to be dynamic by allowing stochastic depths at the transformer encoder. The second step involves creating a derived dataset from the training dataset which contains a pair of image and layer number that provides the highest segmentation quality. Finally, the third step involves training a gating network, whose function is to decide the number of layers to be used given the input image. Extensive experiments demonstrate that ECO-M2F achieves significantly reduced computational complexity compared to established methods while maintaining competitive performance in universal segmentation. Our results highlight ECO-M2F’s ability to dynamically trade-off between performance and efficiency as per requirements, showcasing its adaptability across diverse architectural configurations, and can be applied to models for object detection tasks.
Limitations.